A rectangle's length is seven centimeters more than its width. The rectangle's area is 78 square centimeters. What is its length?
Answer by richard1234(7193)

For some of these area problems it helps to try factoring 78 before doing anything else. Since 78 = 6*13, and 13 = 6+7, we know right away this works. Therefore the length is 13 cm.

Or you can let L be the length, then L-7 is the width (since the length is 7 cm longer). Then L(L-7) = 78. Solve for L.